Yave Number, cm 4000 2000 2500 2000 1500 1200 1200 1100 1000 LLLLLLLLLL 100 Absorbance Wavelength, microns An unknow compound has the molecular formula C5H10O2 has the IR spectrum, and the RMN-^1H show below.

a) Propose a structure that is consistent withthe information provided.

b) Assign all signals in the spectrum of RMN-^1H.

c) Assign at least two bands of the IR spectrum. Note : Remember determine the degrees of unsaturation of the compound.
Answer #1

. Molecular formula is Cato 2 Degrees of unsaturaction/double bond. equivalent is a C+1 -4 = 5+1-10 = 1. a from IR and 14 nmo